each stroke of a vacuum pump removes one third of the air ramianing in a container. wut percent of the original quanity of the air ramains in the container after 10 strokes, to the nearest percentage?
we r workin with geometric sequence

Answer by checkley75(3666)

SINCE EACH STROKE REMOVES 1/3 OF THE AIR THEN THE REMAINING AIR IS 2/3 WHICH IS AGAIN REDUCED BY 1/3 WITH EACH SUCCESSIVE STROBE LEAVING 2/3 OF THE REMAING AIR, ETC. SO AFTER 10 STROKES WE HAVE LEFT:
2/3*2/3*2/3*2/3*2/3*2/3*2/3*2/3*2/3*2/3=1,024/59,049=.0173415=1.7% OF THE AIR IS LEFT.
OR
(2/3)^10=.0173415=1.7%